Imagine an economy that does not have international trade and is initially in equilibrium. In this economy, the marginal propensity to consume is 0.65, and there are no taxes.

Refer to Scenario 10.1. Calculate the value of the spending multiplier for this economy.
Answer #1

Marginal propensity to consume, MPC = 0.65

Spending multiplier = 1/(1-MPC) = 1/(1-0.65) = 1/0.35 = 2.86

The value of the spending multiplier for this economy is 2.86